OpenStudy (anonymous):

subtract: (x+4)/(x-4) - (3x+12)/(x^2-16)

OpenStudy (anonymous):

\[\LARGE \frac{x+4}{x-4}-\frac{3x+12}{x^2-16}\] first you need to find common denominator. Do you know how?

OpenStudy (anonymous):

i did that and ended up with x^2+11x +28

OpenStudy (anonymous):

\[\LARGE {x+4\over x-4}-{3x+12\over x^2-16}=\] \[\LARGE {x+4\over x-4}-{3(x+4)\over (x-4)(x+4)}=\] \[\LARGE {x+4\over x-4}-{3\cancel{(x+4)}\over (x-4)\cancel{(x+4)}}=\] \[\LARGE {x+4\over x-4}-{3\over (x-4)}=\] \[\LARGE {x+4-3\over x-4}={x+1\over x-4}\]